OpenStudy (anonymous):

Find the product of (x + 7)(x − 7). x2 − 14x + 49 x2 − 49 x2 + 14x + 49 x2 + 49 Please help I got the answer x2 + 14x + 49 correct me if I'm wrong ???

OpenStudy (anonymous):

Its very simple my dear friend... we know that (x-a)(x+a)=(x^2-a^2)

OpenStudy (anonymous):

its nothing but a=7 then we get x^2-7^2 = x^2-49 so, B is correct option....

OpenStudy (anonymous):

Thank you so much @chetan552

OpenStudy (anonymous):

You are always welcome.....@NadiaElhaj_1

OpenStudy (anonymous):

FOIL